# The finite groups with three automorphism orbits

### v1

@@ Abstract.
automorphism group; automorphism orbit; exterior squre
@@ 1. Introduction
For a group $N$, the natural action of the automorphism group ${\mathrm{Aut}}(N)$ on $N$ partitions the elements of $N$ into orbits, called the automorphism orbits of $N$. We remark that an automorphism orbit is also called a fusion class. The study of the relations between groups and their automorphism orbits was initiated by G. Higman [16] in 1963 when he studied finite $2$-groups whose involutions form an automorphism orbit, and since then automorphism orbits of groups have received considerable attention of group theorists.
@@
For convenience, a group $N$ is called $r$-orbit if ${\mathrm{Aut}}(N)$ partitions elements of $N$ into $r$ orbits. Since elements in the same automorphism orbit have the same order, the identity group is the only $1$-orbit group, and elementary abelian $p$-groups with $p$ prime are the only $2$-orbit groups. The class of finite $3$-orbit groups contains some interesting families of groups, and the problem of classifying such groups has received considerable attention in the literature. The main purpose of this paper is to classify all finite $3$-orbit groups.
@@
Obviously, a $3$-orbit group has order divisible by at most two primes, and so it is either a $p$-group or a $\{p,q\}$-group, where $p,q$ are distinct primes. It is not hard to see that finite $3$-orbit abelian groups are homocyclic group with exponent $p^{2}$ with $p$ prime. Finite $\{p,q\}$-groups that are $3$-orbit groups have been classified in [10, 12, 20, 24] (also see Example 2.1), which are proved to be certain Frobenius groups. Recently, finite $2$-groups that are $3$-orbit groups are classified independently by Bors and Glasby [4], and the authors [22], with two typical examples being Sylow $2$-subgroups of unitary groups ${\mathrm{SU}}(3,2^{f})$ and Suzuki groups ${\mathrm{Sz}}(2^{2f+1})$ for $f\geqslant 1$ (see Section 2.1). The problem of classifying $3$-orbit groups is therefore reduced to the $p$-groups with odd prime $p$.
@@
As comparison, we would like to mention some interesting researches about automorphism orbits. Studies of $r$-orbit groups for small values of $r>3$ mainly focus on non-solvable groups. Finite simple $r$-orbit groups for $r\leqslant 100$ are determined in [17], and non-solvable $r$-orbit groups for $r\in\{4,5,6\}$ are classified (see [1, 9, 20]). Finite groups with some special automorphism orbits also receive a lot of attention. Shult proved that finite $p$-group for odd prime $p$ such that elements of order $p$ form an automorphism orbit is abelian [27]. In 1976, Gross studied finite $2$-groups whose involutions form an automorphism orbits [14], and the authors complete the classification of such groups recently [22]. Influenced by Gross’ work, finite groups whose elements of the same order lie in the same automorphism orbit, so-called AT-groups, are studied by Zhang [32]. Praeger and the first-named author [21] give a description of finite groups such that elements of the same order are fused or inverse-fused.
@@
Recall that a non-abelian $p$-group $N$ is said to be a special $p$-group if ${\mathbf{Z}}(N)=\Phi(N)=N^{\prime}$ is elementary abelian. In this paper, for a prime power $q=p^{f}$, we denote by $q^{m+n}$ a non-abelian special $p$-group $N$ with $|\Phi(N)|=q^{m}$ and $|N/\Phi(N)|=q^{n}$. We remark that finite non-abelian $3$-orbit $p$-groups are special $p$-groups since it has a unique non-trivial proper characteristic subgroup. For odd prime $p$, finite non-abelian $3$-orbit $p$-groups are special $p$-groups of exponent $p$. It is known that the $n$-generator free group in the variety of groups having exponent $p$ and nilpotency class $2$ is finite, which is denoted by ${\mathcal{H}}_{n,p}$ in this paper (see Definition 2.9). Hence, for odd prime $p$, each non-abelian $n$-generator $3$-orbit $p$-group is isomorphic to a quotient group ${\mathcal{H}}_{n,p}/W$ for some $W<{\mathbf{Z}}({\mathcal{H}}_{n,p})$. To classify $3$-orbit groups, we have to determine all “desired” subgroups $W$. Denote by $\Lambda^{2}_{{\mathrm{F}}}(V)$ the exterior square of vector space $V$ over ${\mathrm{F}}$. We remark that ${\mathbf{Z}}({\mathcal{H}}_{n,p})$ can be “viewed” as the exterior square of ${\mathcal{H}}_{n,p}/{\mathbf{Z}}({\mathcal{H}}_{n,p})$ in some sense (see Lemma 2.10 for more details). The following theorem on representations of finite transitive linear groups plays an important role in the classification.
@@
    $G_{0}:=G^{(\infty)}\cong{\mathrm{SL}}(3,p^{n/3})$ and $\dim\Lambda^{2}_{{\mathbb{F}}_{p}}(V)/W=\dim V=n$.
@@
    $G_{0}:=G^{(\infty)}\cong{\mathrm{Sp}}(2m,p^{n/(2m)})$ and $\Lambda^{2}_{{\mathbb{F}}_{p}}(V)/W$ is a trivial ${\mathbb{F}}_{p}G_{0}$-module with dimension no more than $n/2m$.
Now we present our main theorem on the classification of finite $3$-orbit groups.
@@
(1) ${\mathbb{Z}}_{p}^{n(q-1)}{:}{\mathbb{Z}}_{q}$ ${\mathrm{A}\Gamma\mathrm{L}}(n,p^{q-1})$ Example 2.1 2.2
(2) ${\mathbb{Z}}_{p^{2}}^{n}$ ${\mathrm{GL}}(n,{\mathbb{Z}}/p^{2}{\mathbb{Z}})\cong p^{n^{2}}.{\mathrm{GL}}(n,p)$ 2.3
(3) $A_{2}(n,\theta)\cong 2^{n+n}$ $2^{n^{2}}{:}{\Gamma\mathrm{L}}(1,2^{n})$ $|\theta|\neq 1$ is odd 2.6
(4) ${\mathrm{SU}}(3,2^{n})_{2}\cong 2^{n+2n}$ $2^{2n^{2}}{:}{\Gamma\mathrm{L}}(1,2^{2n})$ 2.6
(5) $P(\epsilon)\cong 2^{3+6}$ $2^{18}{:}({\mathbb{Z}}_{7}{:}{\mathbb{Z}}_{9})$ 2.6
(6) $A_{p}(n,\theta)\cong p^{n+n}$ $p^{n^{2}}{:}{\Gamma\mathrm{L}}(3,p^{n/3})$ $p$ is odd and $|\theta|=3$ 4.3
(7) $q^{1+2m}_{+}\cong p^{n+2mn}$ $p^{2mn^{2}}{:}({\mathrm{Sp}}(2m,q){:}{\Gamma\mathrm{L}}(1,q))$ $q=p^{n}$ and $p$ is odd 4.4
(8) $q^{1+2m}_{+}/U\cong p^{n_{0}+2mn}$ $p^{2mn_{0}n}{:}({\mathrm{Sp}}(2m,q){:}{\Gamma\mathrm{L}}(1,q)_{U})$ Example 5.3 5.4
In Table 1, groups $A_{p}(n,\theta)$’s are known as Suzuki $p$-group of type A (see Example 2.4); the group $P(\epsilon)$ was first given by Dornhoff [12] (see Theorem 2.6); and $q^{1+2m}_{+}$ is a generalization of $p^{1+2m}_{+}$, the extraspecial $p$-group of exponent $p$ with odd prime $p$. We now explain how to generalize $p^{1+2m}_{+}$ to $q^{1+2m}_{+}$.
Let $q=p^{n}$ for an odd prime $p$. For convenience, we denote a Sylow $p$-subgroup of a group $X$ by $X_{p}$. It is known that $p^{1+2}_{+}\cong{\mathrm{SL}}(3,p)_{p}$ and $p^{1+2m}_{+}$ is the central product of $m$-copies of $p^{1+2}_{+}$. A natural way to generalize such groups is defining $q^{1+2}_{+}$ to be ${\mathrm{SL}}(3,q)_{p}$ and defining $q^{1+2m}_{+}$ to be the central product of $m$-copies of $q^{1+2}_{+}$ (also see Definition 2.17). A classical result of Winter [31] says that ${\mathrm{Out}}(p^{1+2m}_{+})\cong\mathrm{CSp}(2m,p)\cong{\mathrm{Sp}}(2m,p){:}{\mathbb{Z}}_{p-1}$. In Lemma 4.4, we prove that ${\mathrm{Aut}}(q^{1+2m}_{+})\cong p^{2mn^{2}}{:}{\mathrm{C}\Gamma\mathrm{Sp}}(2m,q)$, which shows some similarities between $p^{1+2m}_{+}$ and $q^{1+2m}_{+}$. Thus, we name the group $q^{1+2m}_{+}$ the extraspecial $q$-group.
@@
Let $N:=q^{1+2m}_{+}$, and let $H:={\mathrm{Aut}}(N)^{{\mathbf{Z}}(N)}\cong{\Gamma\mathrm{L}}(1,q)$. Then $N/U$ is a $3$-orbit group with $U<{\mathbf{Z}}(N)$ if and only if $H_{U}$ acts transitively on non-identity elements of ${\mathbf{Z}}(N)/U\cong{\mathbf{Z}}(N/U)$.
Additionally, we obtain some isomorphisms among some $p$-groups related to extraspecial $q$-groups, which is independently interesting. We remark that an ${\mathbb{F}}_{p}$-subspace $U$ of $V={\mathbb{F}}_{p^{n}}^{+}$ is called a subfield hyperplane with respect to subfield ${\mathbb{F}}_{p^{d}}$ for $d\mid n$ if and only if $U$ is a hyperplane of $V$ as ${\mathbb{F}}_{p^{d}}$-spaces (also see Definition 5.1).
@@
A transitive permutation group $G$ on a set ${\Omega}$ is said to be of rank $r$ if $G$ has exactly $r$ orbits on ${\Omega}\times{\Omega}$. The study of rank $3$ groups dates back to work of G. Higman [15] in 1964. Note that the holomorph ${\mathrm{Hol}}(N)=N{:}{\mathrm{Aut}}(N)$ of a group $N$ has rank $r$ if and only if $N$ is an $r$-group. Hence, our classification of $3$-orbit groups provides new families of rank $3$ groups.
@@
The holomorph ${\mathrm{Hol}}(N)\cong N{:}{\mathrm{Aut}}(N)$ is a finite rank $3$ group if and only if $N$ is one of groups in Table 1.
@@ Acknowledgments
This work was supported by NNSFC grant no. 11931005. The authors thank to Eamonn A. O’Brien for his advice on Magma.
@@ 2.1. Some known finite $3$-orbit groups
$N:=P{:}Q\cong{\mathbb{Z}}_{p}^{n(q-1)}{:}{\mathbb{Z}}_{q}$
@@
We make the following observations on the group $N:=P{:}Q$.
@@
    Viewing $P\cong{\mathbb{Z}}_{p}^{n(q-1)}$ as a vector space, $Q$ acting on $P$ is a scalar multiplication by definition, and so $P$ is a homogeneous ${\mathbb{F}}_{p}Q$-module. Thus, $N$ is a homogeneous Frobenius group, and ${\mathrm{Aut}}(N)\cong X={\mathrm{A}\Gamma\mathrm{L}}(n,p^{q-1})$, by [29, Theorem 1.1].
@@
    Elements of $N$ of order $p$ are conjugate under ${\Gamma\mathrm{L}}(n,p^{q-1})$. Sylow $q$-subgroups of $N$ are conjugate. The field automorphism $\phi$ of ${\mathbb{F}}_{p^{q-1}}$ acts on $Q$ faithfully, and so non-identity elements of $Q$ are conjugate under $\langle\phi\rangle$. Thus, elements in $N$ of order $q$ are conjugate under ${\mathrm{Aut}}(N)$, and $N=P{:}Q$ is a $3$-orbit group.
@@
A finite group $3$-orbit group $N$ is not a $p$-group if and only if $N\cong{\mathbb{Z}}_{p}^{n(q-1)}{:}{\mathbb{Z}}_{q}$, as described in Example 2.1. In this case, ${\mathrm{Aut}}(N)\cong{\mathrm{A}\Gamma\mathrm{L}}(n,p^{q-1})$.
It is known and easily shown that each homocyclic group ${\mathbb{Z}}_{p^{2}}^{n}$ is a $3$-orbit group, and
@@
A non-abelian $2$-group with more than one involutions is called a Suzuki $2$-group if it has a cyclic group of automorphisms acting transitively on involutions. D. Higman classified all Suzuki $2$-group and divided them into four types A-D in [16]. The definition of each type of Suzuki $2$-group can be extended to odd primes, which are called Suzuki $p$-groups. The following is the definition of Suzuki $p$-group of type A, also see [2, Section 46].
@@
$A_{p}(n,\theta):=\langle M(a,b,\theta)\mid a,b\in{\mathbb{F}}_{p^{n}}\rangle,$
@@
We remark that groups $A_{2}(n,\theta)$ with $|\theta|\neq 1$ of odd order form the class of Suzuki $2$-groups of type A given by D. Higman. We collect some properties of $A_{p}(n,\theta)$ blow.
@@
Let $N:=A_{p}(n,\theta)$. Then the following statements hold.
@@
    [22, Lemma 4.2] When $p=2$ and $|\theta|>1$ is odd, $N$ is a $3$-orbit $2$-group with ${\mathrm{Aut}}(N)\cong 2^{n^{2}}{:}{\Gamma\mathrm{L}}(1,2^{n})$.
@@
    Let $\lambda$ be a generator of ${\mathbb{F}}_{p^{n}}$, and let $D=\mathrm{diag}(\lambda^{-1},1,\lambda^{\theta})$. Then $D$ induces an an automorphism $\xi$ of $N$ such that
@@
We remark that $A_{2}(2f+1,\theta)$ for $x^{\theta}=x^{2^{f+1}}$ is isomorphic to a Sylow $2$-subgroup of Suzuki group ${\mathrm{Sz}}(2^{2f+1})$, see [8, Proposition 13.4(vi)]. Hence, Sylow $2$-subgroups of ${\mathrm{Sz}}(2^{2f+1})$ are $3$-orbit $2$-groups.
Finite $3$-orbit $2$-groups are classified independently by Bors and Glasby [4], and the authors [22]. We record the results blow.

### v2

@@ Abstract.
automorphism group; automorphism orbit; exterior square
@@ 1. Introduction
For a group $N$, the natural action of the automorphism group ${\mathrm{Aut}}(N)$ on $N$ partitions the elements of $N$ into orbits, called the automorphism orbits of $N$. We remark that an automorphism orbit is also called a fusion class. The study of the relations between groups and their automorphism orbits was initiated by G. Higman [19] in 1963 when he studied finite $2$-groups whose involutions form an automorphism orbit, and since then automorphism orbits of groups have received considerable attention from group theorists.
@@
For convenience, a group $N$ is called $r$-orbit if ${\mathrm{Aut}}(N)$ partitions the elements of $N$ into $r$ orbits. Since elements in the same automorphism orbit have the same order, the trivial group is the only $1$-orbit group, and elementary abelian $p$-groups with $p$ prime are the only finite $2$-orbit groups. It is worth mentioning that there are examples of non-abelian infinite $2$-orbit groups, see [30, Corollary 1.3]. The class of finite $3$-orbit groups contains some interesting families of groups, and the problem of classifying such groups has received considerable attention in the literature. The main purpose of this paper is to classify all finite $3$-orbit groups.
@@
Obviously, a $3$-orbit group has order divisible by at most two primes, and so it is either a $p$-group or a $\{p,q\}$-group, where $p,q$ are distinct primes. It is not hard to see that finite $3$-orbit abelian groups are homocyclic groups with exponent $p^{2}$ with $p$ prime. Finite $\{p,q\}$-groups that are $3$-orbit groups have been classified in [12, 14, 23, 27] (also see Example 2.1), which are proved to be certain Frobenius groups. Mäurer and Stroppel [27] study groups $3$-orbit groups under additional structural assumptions. Recently, finite $2$-groups that are $3$-orbit groups have been classified independently by Bors and Glasby [6], and the authors [25], with two typical examples being Sylow $2$-subgroups of unitary groups ${\mathrm{SU}}(3,2^{f})$ and Suzuki groups ${\mathrm{Sz}}(2^{2f+1})$ for $f\geqslant 1$ (see Section 2.1). The problem of classifying $3$-orbit groups is therefore reduced to the the case of $p$-groups with $p$ an odd prime.
@@
As comparison, we would like to mention some interesting research about automorphism orbits. Studies of $r$-orbit groups for small values of $r>3$ mainly focus on non-solvable groups. Specifically, ${\mathrm{A}}_{5}$ is the only $4$-orbit non-solvable group in [23, Theorem 3], and ${\mathrm{PSL}}(2,p)$ for $p\in\{7,8,9\}$ are the only $5$-orbit non-abelian simple groups [34]. The finite simple $r$-orbit groups for $r\leqslant 100$ are determined in [20], and the non-solvable $r$-orbit groups for $r=5,6$ are classified in [2, 11]. Finite groups with some special automorphism orbits also receive a lot of attention. Shult proved that finite $p$-group for odd prime $p$ in which the elements of order $p$ form an automorphism orbit is abelian [32]. In 1976, Gross studied finite $2$-groups whose involutions form an automorphism orbit [16], and the authors have recently completed the classification of such groups recently [25]. Influenced by Gross’ work, finite groups whose elements of the same order lie in the same automorphism orbit, so-called AT-groups, were studied by Zhang [38]. Praeger and the first-named author [24] gave a description of finite groups in which elements of the same order are fused or inverse-fused. Recently, Bors, Giudici and Praeger studied the relation between the number of automorphism orbits and the size of the spectrum of a non-solvable group [5].
@@
Recall that a non-abelian $p$-group $N$ is said to be a special $p$-group if ${\mathbf{Z}}(N)=\Phi(N)=N^{\prime}$ is elementary abelian. In this paper, for a prime power $q=p^{f}$, we denote by $q^{m+n}$ a non-abelian special $p$-group $N$ with $|\Phi(N)|=q^{m}$ and $|N/\Phi(N)|=q^{n}$. We remark that finite non-abelian $3$-orbit $p$-groups are special $p$-groups since each of them has a unique non-trivial proper characteristic subgroup. For odd prime $p$, finite non-abelian $3$-orbit $p$-groups are special $p$-groups of exponent $p$. It is known that the $n$-generator free group in the variety of groups having exponent $p$ and nilpotency class $2$ is finite, which is denoted by ${\mathcal{H}}_{n,p}$ in this paper (see Definition 2.9). Hence, for odd prime $p$, each non-abelian $n$-generator $3$-orbit $p$-group is isomorphic to a quotient group ${\mathcal{H}}_{n,p}/W$ for some $W<{\mathbf{Z}}({\mathcal{H}}_{n,p})$. To classify $3$-orbit groups, we have to determine all “desired” subgroups $W$. Denote by $\Lambda^{2}_{{\mathrm{F}}}(V)$ the exterior square of vector space $V$ over ${\mathrm{F}}$. We remark that ${\mathbf{Z}}({\mathcal{H}}_{n,p})$ can be “viewed” as the exterior square of ${\mathcal{H}}_{n,p}/{\mathbf{Z}}({\mathcal{H}}_{n,p})$ in some sense (see Lemma 2.10 for more details). The following theorem, an analog of [25, Theorem 1.2] which is for $p=2$, describes certain representations of finite transitive linear groups of characteristic $p$.
@@
    $G^{(\infty)}\cong{\mathrm{SL}}(3,p^{n/3})$ and $\dim\Lambda^{2}_{{\mathbb{F}}_{p}}(V)/W=\dim V=n$.
@@
    $G^{(\infty)}\cong{\mathrm{Sp}}(2m,p^{n/(2m)})$ and $\Lambda^{2}_{{\mathbb{F}}_{p}}(V)/W$ is a trivial ${\mathbb{F}}_{p}G^{(\infty)}$-module with dimension at most $n/(2m)$.
This result plays an important role in proving the second main theorem of the paper, which presents a classification of finite $3$-orbit groups, stated as follows.
@@
1 ${\mathbb{Z}}_{p}^{n(q-1)}{:}{\mathbb{Z}}_{q}$ ${\mathrm{A}\Gamma\mathrm{L}}(n,p^{q-1})$ Example 2.1 2.2
2 ${\mathbb{Z}}_{p^{2}}^{n}$ ${\mathrm{GL}}(n,{\mathbb{Z}}/p^{2}{\mathbb{Z}})\cong p^{n^{2}}.{\mathrm{GL}}(n,p)$ 2.3
3 $A_{2}(n,\theta)\cong 2^{n+n}$ $2^{n^{2}}{:}{\Gamma\mathrm{L}}(1,2^{n})$ $|\theta|\neq 1$ is odd 2.6
4 ${\mathrm{SU}}(3,2^{n})_{2}\cong 2^{n+2n}$ $2^{2n^{2}}{:}{\Gamma\mathrm{L}}(1,2^{2n})$ 2.6
5 $P(\epsilon)\cong 2^{3+6}$ $2^{18}{:}({\mathbb{Z}}_{7}{:}{\mathbb{Z}}_{9})$ 2.6
6 $A_{p}(n,\theta)\cong p^{n+n}$ $p^{n^{2}}{:}{\Gamma\mathrm{L}}(3,p^{n/3})$ $p$ is odd and $|\theta|=3$ 4.3
7 $q^{1+2m}_{+}\cong p^{n+2mn}$ $p^{2mn^{2}}{:}({\mathrm{Sp}}(2m,q){:}{\Gamma\mathrm{L}}(1,q))$ $q=p^{n}$ and $p$ is odd 4.4
8 $q^{1+2m}_{+}/U\cong p^{n_{0}+2mn}$ $p^{2mn_{0}n}{:}({\mathrm{Sp}}(2m,q){:}{\Gamma\mathrm{L}}(1,q)_{U})$ Example 5.3 5.4
In Table 1, the groups $A_{p}(n,\theta)$ are known as Suzuki $p$-groups of type A (see Example 2.4); the group $P(\epsilon)$ was first given by Dornhoff [14] (see Theorem 2.6); and $q^{1+2m}_{+}$ is a generalization of $p^{1+2m}_{+}$, the extraspecial $p$-groups of exponent $p$ with odd prime $p$. We now explain how to generalize $p^{1+2m}_{+}$ to $q^{1+2m}_{+}$.
Let $q=p^{n}$ for an odd prime $p$. For convenience, we denote a Sylow $p$-subgroup of a group $X$ by $X_{p}$. It is known that $p^{1+2}_{+}\cong{\mathrm{SL}}(3,p)_{p}$ and $p^{1+2m}_{+}$ is the central product of $m$ copies of $p^{1+2}_{+}$. A natural way to generalize such groups is defining $q^{1+2}_{+}$ to be ${\mathrm{SL}}(3,q)_{p}$ and defining $q^{1+2m}_{+}$ to be the central product of $m$ copies of $q^{1+2}_{+}$ (also see Definition 2.17). A classical result of Winter [37] says that ${\mathrm{Out}}(p^{1+2m}_{+})\cong\mathrm{CSp}(2m,p)\cong{\mathrm{Sp}}(2m,p){:}{\mathbb{Z}}_{p-1}$. In Lemma 4.4, we prove that ${\mathrm{Aut}}(q^{1+2m}_{+})\cong p^{2mn^{2}}{:}{\mathrm{C}\Gamma\mathrm{Sp}}(2m,q)$, which shows some similarities between $p^{1+2m}_{+}$ and $q^{1+2m}_{+}$. Thus, we name the group $q^{1+2m}_{+}$ the extraspecial $q$-group.
@@
Let $N=q^{1+2m}_{+}$, and let $H={\mathrm{Aut}}(N)^{{\mathbf{Z}}(N)}\cong{\Gamma\mathrm{L}}(1,q)$. Then $N/U$ is a $3$-orbit group with $U<{\mathbf{Z}}(N)$ if and only if $H_{U}$ acts transitively on the non-identity elements of ${\mathbf{Z}}(N)/U\cong{\mathbf{Z}}(N/U)$.
Moreover, we give some interesting properties of extraspecial $q$-groups. We remark that an ${\mathbb{F}}_{p}$-subspace $U$ of $V={\mathbb{F}}_{p^{n}}^{+}$ is called a subfield hyperplane with respect to subfield ${\mathbb{F}}_{p^{d}}$ for $d\mid n$ if and only if $U$ is a hyperplane of $V$ as ${\mathbb{F}}_{p^{d}}$-spaces (also see Definition 5.1).
@@
Let $N$ be a non-abelian $3$-orbit $p$-group with center ${\mathbb{Z}}_{p}$ and exponent $p$. Theorems B and C (i) imply that $N$ is an extraspecial $p$-group $p^{1+2m}_{+}$ if $N$ is finite. Mäurer and Stroppel studied the structure of such infinite groups [27, Theorem 6.4].
A transitive permutation group $G$ on a set $\Omega$ is said to be of rank $r$ if $G$ has exactly $r$ orbits on $\Omega\times\Omega$. The study of rank $3$ groups dates back to work of G. Higman [18] in 1964. Note that the holomorph ${\mathrm{Hol}}(N)=N{:}{\mathrm{Aut}}(N)$ of a group $N$ has rank $r$ if and only if $N$ is an $r$-orbit group. Recently, Huang and the authors established a reduction theorem towards a classification of the finite rank $3$ permutation groups [17, Corollary 3], of which case (iii) comprises the groups with regular normal subgroups. Hence, our classification of $3$-orbit groups provides the “maximal” groups in this case.
@@
The holomorph ${\mathrm{Hol}}(N)\cong N{:}{\mathrm{Aut}}(N)$ is a finite rank $3$ permutation group if and only if $N$ is one of the groups in Table 1.
@@
We will introduce some known finite $3$-orbit groups in Section 2.1, for example, $\{p,q\}$-groups, abelian groups and $2$-groups. Then some basic properties of special $p$-groups of exponent $p$ will be given in Section 2.2. A new family of $3$-orbit $p$-group, the extraspecial $q$-groups, will be constructed in Section 2.3.
The proof of Theorem A will be given in Section 3, which is a generalization of [25, Theorem 1.1]. With the aid of this theorem, we obtain the full automorphism group of the groups in rows 1-7 of Table 1 in Section 4.
We will discuss some properties of quotients of extraspecial $q$-groups in Section 5, and then complete proofs of Theorems B and C.
@@ Acknowledgments
This work was supported by NNSFC grant no. 11931005. The authors thank Eamonn A. O’Brien for his advice on Magma computations, and are grateful to the anonymous referees for valuable comments and suggestions, which have helped improve the paper.
@@ 2.1. Some known finite $3$-orbit groups
$P{:}Q\cong{\mathbb{Z}}_{p}^{n(q-1)}{:}{\mathbb{Z}}_{q}$
@@
Let $N=P{:}Q$ as defined in Example 2.1. We make the following observations on $N$.
@@
    Viewing $P\cong{\mathbb{Z}}_{p}^{n(q-1)}$ as a vector space, $Q$ acts on $P$ via scalar multiplications by definition, and so $P$ is a homogeneous ${\mathbb{F}}_{p}Q$-module. Thus, $N$ is a homogeneous Frobenius group, see [23, Theorem 2]. Then, by [35, Theorem 1.1], we can easily obtain that ${\mathrm{Aut}}(N)\cong X={\mathrm{A}\Gamma\mathrm{L}}(n,p^{q-1})$.
@@
    Elements in $N$ of order $p$ are conjugate under ${\Gamma\mathrm{L}}(n,p^{q-1})$. Sylow $q$-subgroups of $N$ are conjugate. The field automorphism $\phi$ of ${\mathbb{F}}_{p^{q-1}}$ acts on $Q$ faithfully, and so non-identity elements of $Q$ are conjugate under $\langle\phi\rangle$. Thus, elements in $N$ of order $q$ are conjugate under ${\mathrm{Aut}}(N)$, and $N=P{:}Q$ is a $3$-orbit group.
@@
A finite $3$-orbit group $N$ is not a $p$-group if and only if $N\cong{\mathbb{Z}}_{p}^{n(q-1)}{:}{\mathbb{Z}}_{q}$, as described in Example 2.1. In this case, ${\mathrm{Aut}}(N)\cong{\mathrm{A}\Gamma\mathrm{L}}(n,p^{q-1})$.
Obviously, each homocyclic group ${\mathbb{Z}}_{p^{2}}^{n}$ is a $3$-orbit group, and
@@
We note that Schwachhöfer and Stroppel [31] studied the structure of $k$-orbit abelian groups (especially for infinite groups) for finite $k$.
A non-abelian $2$-group with more than one involution is called a Suzuki $2$-group if it has a cyclic group of automorphisms acting transitively on involutions. D. Higman classified all Suzuki $2$-groups and divided them into four types A-D in [19]. The definition of each type of Suzuki $2$-groups can be extended to odd primes, which are called Suzuki $p$-groups. The following is the definition of Suzuki $p$-group of type A, see [3, Section 46].
@@
$A_{p}(n,\theta)=\langle M(a,b,\theta)\mid a,b\in{\mathbb{F}}_{p^{n}}\rangle,$
@@
We remark that groups $A_{2}(n,\theta)$ with $|\theta|\neq 1$ of odd order form the class of Suzuki $2$-groups of type A given by D. Higman, which are $3$-orbit groups. The following proposition can be easily verified, where the first part was given in [25, Lemma 4.2].
@@
Let $N=A_{p}(n,\theta)$. Then the following statements hold.
@@
    When $p=2$ and $|\theta|>1$ is odd, $N$ is a $3$-orbit $2$-group with ${\mathrm{Aut}}(N)\cong 2^{n^{2}}{:}{\Gamma\mathrm{L}}(1,2^{n})$.
@@
    Let $\lambda$ be a generator of ${\mathbb{F}}_{p^{n}}$, and let $D=\mathrm{diag}(\lambda^{-1},1,\lambda^{\theta})$. Then $D$ induces an automorphism $\xi$ of $N$ such that
@@
We remark that $A_{2}(2f+1,\theta)$ for $x^{\theta}=x^{2^{f+1}}$ is isomorphic to a Sylow $2$-subgroup of Suzuki group ${\mathrm{Sz}}(2^{2f+1})$, see [9, Proposition 13.4 (vi)]. Hence, Sylow $2$-subgroups of ${\mathrm{Sz}}(2^{2f+1})$ are $3$-orbit $2$-groups.
Finite $3$-orbit $2$-groups are classified independently by Bors and Glasby [6], and the authors [25]. We record the results below.
@@
We remark that groups in parts (ii) and (iii) are Suzuki $2$-groups of type B in Higman’s notation, see [14, page 706] and [6, Theorem 1.2].